# The LLM Data Auditor: A Metric-oriented Survey on Quality and Trustworthiness in Evaluating Synthetic Data

### v1

@@ 6.2.2 Video-Text
To evaluate dynamic variety, we measure motion magnitude using a motion score based on optical flow(Teed and Deng, 2020). This score is calculated using the following equation:
@@ 6.5 Usages
Similarly, LLaVAR-2 integrates human-annotated captions with filtered instructions generated by GPT-4o to curate a high-quality and text-centric supervised fine-tuning dataset (Zhou et al., 2024). For the purpose of targeting instance-level grounding, Inst-IT synthesizes explicit visual-prompted instructions for continuous supervised fine-tuning. This method enhances instance-grounded understanding and demonstrates effective transfer to generic benchmarks (Peng et al., 2024). Concurrently, dense and high-fidelity synthetic captions have proven effective for supervision. For example, ShareGPT4V and ShareGPT4Video demonstrate that large-scale synthetic captioning directly benefits both supervised finetuning and unified large vision-language model pre-training (Chen et al., 2023b; 2024a).
@@ 7.1 Generation Methods
In urban traffic simulation settings, LLMs parse free-form requests into structured keywords such as a dictionary of scenario fields. These keywords then drive the generation of scriptable configurations and parameters for urban mobility testing (Li et al., 2024c). Multimodal pipelines extend this concept to rare and difficult situations by generating realistic corner cases along with runnable tests that target the long tail of driving scenarios (Lu et al., 2024). Other complementary pipelines convert operational design domain descriptions into scripts compatible with ScenarioRunner for simulation-based testing (Danso and Büker, 2025).
@@
In the area of embodied artificial intelligence, long-horizon planners such as L3M+P generate PDDL problems and maintain world-state graphs that can be instantiated into many concrete tasks and goals (Agarwal et al., 2025b). Safety-focused planners such as SELP translate natural language tasks into temporal logic specifications and use equivalence voting to improve the robustness and consistency of the mapping from natural language to logic. Following this, constrained planning enforces the resulting safety constraints during execution (Wu et al., 2025). Self-correcting frameworks such as the T3 Planner verify plans against spatio-temporal logic and repair them when needed. This process produces verified plan and trajectory traces that can be recorded as supervision for evaluation or learning under explicit constraints (Li and Zhao, 2025). Large-scale multi-agent benchmarks such as PARTNR also belong to this category. In these benchmarks, LLMs help design and decompose collaboration tasks, ground them into simulators, and export them as standard task collections together with planner baselines (Chang et al., 2024).
@@
In embodied artificial intelligence, LLMs can also augment demonstration and trajectory data more directly. For instance, LLM Trainer performs offline demonstration annotation and online keypose retargeting to adapt demonstrations to new scenes and generate additional imitation trajectories with minimal human input (George and Farimani, 2025). Frameworks such as ELLMER integrate high-level language-driven task decomposition with low-level execution using vision or force feedback and retrieval-augmented code examples to solve long-horizon tasks. The resulting executions naturally produce rich interaction traces that can be recorded for analysis (Mon-Williams and others, 2025). Program-structured approaches such as Instruct2Act and ProgPrompt leverage executable or program-like representations to ground instructions in available actions as well as perception and planning loops in some systems. This enables a more modular analysis of which components are necessary (Huang et al., 2023; Singh et al., 2022). When execution logs contain both actions and rich observations, we treat them as control and decision data in this survey if their main use is to study or improve behavior rather than perception.
@@ 7.3 Trustworthy Metrics for Agent Data
Safety measures whether trajectories, scenarios, and plans generated by LLMs avoid hazardous states and rule-breaking behaviors and satisfy explicit safety constraints, independent of task completion. In practice, safety is often assessed through infraction and collision metrics defined by simulators in driving benchmarks, such as suites based on CARLA like SafeBench. It is also evaluated through the satisfaction of formal specifications in pipelines for constraint-enforced planning and motion planning (Xu et al., 2022; Wu et al., 2025; Li and Zhao, 2025). Systems for text-to-scenario generation further emphasize the production of standardized scenario files and simulator testing reports that include monitored evaluation indicators (Cai et al., 2025).
@@
where $V_{i}^{(t)}$ is the number of infractions of type $t$ in episode $i$. This style of reporting is also used in CARLA Leaderboard results where multiple infraction types are shown as infractions per kilometer (CARLA Autonomous Driving Leaderboard, 2025a). Fine-grained violation counts for each specific type additionally enable a more detailed diagnosis than aggregate success or return (Cai et al., 2025; Xu et al., 2022).
@@
where $v_{\min}(t)$ is a context-dependent minimum-speed requirement such as a speed determined by nearby traffic. The CARLA Leaderboard explicitly penalizes the failure to maintain minimum speed as part of its evaluation criteria (CARLA Autonomous Driving Leaderboard, 2025b). More generally, one may also track speed-limit compliance within specific bounds when such information is available in the simulator or map.
@@
The SELP framework improves the safety rate by mapping natural language to Linear Temporal Logic and enforcing constrained decoding, while the T3 Planner utilizes Signal Temporal Logic verification in the loop to increase the satisfaction rates of motion plans (Wu et al., 2025; Li and Zhao, 2025).
@@
Lower values for the minimum Time-to-Collision indicate a higher risk of collision (Ward et al., 2015; Sharath and Mehran, 2021).
@@ 7.4 Evaluation Practice Gap
T3 Planner (Li and Zhao, 2025) $\checkmark$ $\triangle$ $\times$ $\checkmark$ $\checkmark$
@@
Diagnostic Safety. Concerning the diagnosis of safety issues, current reporting relies heavily on broad benchmark metrics such as those related to collisions in evaluations using the CARLA simulator. It also relies on the implicit satisfaction of specifications such as adherence to linear temporal logic or signal temporal logic. These methods are used instead of detailed and specific diagnostics for different types of violations. In our representative set, the generation of traffic scenes reports signals related to collisions. Meanwhile, the SELP(Wu et al., 2025) and T3 Planner (Li and Zhao, 2025) methods mainly focus on the satisfaction of formal constraints. In contrast, other approaches mainly report success at the task level. This includes the completion of tasks or the resolution of violations. However, they do not provide an independent breakdown of safety violations for each specific type. As a result, although current evaluations can show whether safety constraints are met, they provide limited diagnostic detail for identifying and explaining specific modes of safety failure.
@@ 7.5 Usages
Beyond static enforcement, this data also drives self-correction. Systems often pair planners based on LLMs with logic-based verifiers that iteratively reject and repair unsafe actions (Li and Zhao, 2025). This iterative loop can be logged to form trial-and-error traces consisting of failed plans, verifier diagnostics such as constraint violations or robustness scores, and corrected revisions. These traces are useful for evaluation and may be repurposed for future training.

### v2

@@ 6.2.2 Video-Text
To evaluate dynamic variety, we measure motion magnitude using a motion score based on optical flow(Teed & Deng, 2020). This score is calculated using the following equation:
@@ 6.5 Usages
Similarly, LLaVAR-2 integrates human-annotated captions with filtered instructions generated by GPT-4o to curate a high-quality and text-centric supervised fine-tuning dataset (Zhou et al., 2024). For the purpose of targeting instance-level grounding, Inst-IT synthesizes explicit visual-prompted instructions for continuous supervised fine-tuning. This method enhances instance-grounded understanding and demonstrates effective transfer to generic benchmarks (Peng et al., 2024). Concurrently, dense and high-fidelity synthetic captions have proven effective for supervision. For example, ShareGPT4V and ShareGPT4Video demonstrate that large-scale synthetic captioning directly benefits both supervised finetuning and unified large vision-language model pre-training (Chen et al., 2023b, 2024a).
@@ 7.1 Generation Methods
In urban traffic simulation settings, LLMs parse free-form requests into structured keywords such as a dictionary of scenario fields. These keywords then drive the generation of scriptable configurations and parameters for urban mobility testing (Li et al., 2024c). Multimodal pipelines extend this concept to rare and difficult situations by generating realistic corner cases along with runnable tests that target the long tail of driving scenarios (Lu et al., 2024). Other complementary pipelines convert operational design domain descriptions into scripts compatible with ScenarioRunner for simulation-based testing (Danso & Büker, 2025).
@@
In the area of embodied artificial intelligence, long-horizon planners such as L3M+P generate PDDL problems and maintain world-state graphs that can be instantiated into many concrete tasks and goals (Agarwal et al., 2025b). Safety-focused planners such as SELP translate natural language tasks into temporal logic specifications and use equivalence voting to improve the robustness and consistency of the mapping from natural language to logic. Following this, constrained planning enforces the resulting safety constraints during execution (Wu et al., 2025). Self-correcting frameworks such as the T3 Planner verify plans against spatio-temporal logic and repair them when needed. This process produces verified plan and trajectory traces that can be recorded as supervision for evaluation or learning under explicit constraints (Li & Zhao, 2025). Large-scale multi-agent benchmarks such as PARTNR also belong to this category. In these benchmarks, LLMs help design and decompose collaboration tasks, ground them into simulators, and export them as standard task collections together with planner baselines (Chang et al., 2024).
@@
In embodied artificial intelligence, LLMs can also augment demonstration and trajectory data more directly. For instance, LLM Trainer performs offline demonstration annotation and online keypose retargeting to adapt demonstrations to new scenes and generate additional imitation trajectories with minimal human input (George & Farimani, 2025). Frameworks such as ELLMER integrate high-level language-driven task decomposition with low-level execution using vision or force feedback and retrieval-augmented code examples to solve long-horizon tasks. The resulting executions naturally produce rich interaction traces that can be recorded for analysis (Mon-Williams et al., 2025). Program-structured approaches such as Instruct2Act and ProgPrompt leverage executable or program-like representations to ground instructions in available actions as well as perception and planning loops in some systems. This enables a more modular analysis of which components are necessary (Huang et al., 2023; Singh et al., 2022). When execution logs contain both actions and rich observations, we treat them as control and decision data in this survey if their main use is to study or improve behavior rather than perception.
@@ 7.3 Trustworthy Metrics for Agent Data
Safety measures whether trajectories, scenarios, and plans generated by LLMs avoid hazardous states and rule-breaking behaviors and satisfy explicit safety constraints, independent of task completion. In practice, safety is often assessed through infraction and collision metrics defined by simulators in driving benchmarks, such as suites based on CARLA like SafeBench. It is also evaluated through the satisfaction of formal specifications in pipelines for constraint-enforced planning and motion planning (Xu et al., 2022; Wu et al., 2025; Li & Zhao, 2025). Systems for text-to-scenario generation further emphasize the production of standardized scenario files and simulator testing reports that include monitored evaluation indicators (Cai et al., 2025).
@@
where $V_{i}^{(t)}$ is the number of infractions of type $t$ in episode $i$. This style of reporting is also used in CARLA Leaderboard results where multiple infraction types are shown as infractions per kilometer (CARLA Autonomous Driving Leaderboard, 2025b). Fine-grained violation counts for each specific type additionally enable a more detailed diagnosis than aggregate success or return (Cai et al., 2025; Xu et al., 2022).
@@
where $v_{\min}(t)$ is a context-dependent minimum-speed requirement such as a speed determined by nearby traffic. The CARLA Leaderboard explicitly penalizes the failure to maintain minimum speed as part of its evaluation criteria (CARLA Autonomous Driving Leaderboard, 2025a). More generally, one may also track speed-limit compliance within specific bounds when such information is available in the simulator or map.
@@
The SELP framework improves the safety rate by mapping natural language to Linear Temporal Logic and enforcing constrained decoding, while the T3 Planner utilizes Signal Temporal Logic verification in the loop to increase the satisfaction rates of motion plans (Wu et al., 2025; Li & Zhao, 2025).
@@
Lower values for the minimum Time-to-Collision indicate a higher risk of collision (Ward et al., 2015; Sharath & Mehran, 2021).
@@ 7.4 Evaluation Practice Gap
T3 Planner (Li & Zhao, 2025) $\checkmark$ $\triangle$ $\times$ $\checkmark$ $\checkmark$
@@
Diagnostic Safety. Concerning the diagnosis of safety issues, current reporting relies heavily on broad benchmark metrics such as those related to collisions in evaluations using the CARLA simulator. It also relies on the implicit satisfaction of specifications such as adherence to linear temporal logic or signal temporal logic. These methods are used instead of detailed and specific diagnostics for different types of violations. In our representative set, the generation of traffic scenes reports signals related to collisions. Meanwhile, the SELP(Wu et al., 2025) and T3 Planner (Li & Zhao, 2025) methods mainly focus on the satisfaction of formal constraints. In contrast, other approaches mainly report success at the task level. This includes the completion of tasks or the resolution of violations. However, they do not provide an independent breakdown of safety violations for each specific type. As a result, although current evaluations can show whether safety constraints are met, they provide limited diagnostic detail for identifying and explaining specific modes of safety failure.
@@ 7.5 Usages
Beyond static enforcement, this data also drives self-correction. Systems often pair planners based on LLMs with logic-based verifiers that iteratively reject and repair unsafe actions (Li & Zhao, 2025). This iterative loop can be logged to form trial-and-error traces consisting of failed plans, verifier diagnostics such as constraint violations or robustness scores, and corrected revisions. These traces are useful for evaluation and may be repurposed for future training.